Find three consecutive odd integers whose sum is -21.
Let -x=1st #
Let -x+-2=2nd #
Let -x + -2+3rd #
-x+-x+-2+-x+-2=-21
-3x+-4+=-21
As the answer is -9 I am doing something wrong, but what?

Answer by mananth(16949)

let the first odd integer be x
the other will be x+2
the third will be x+4
..
their sum = -21
x+x+2+x+4=-21
3x+6=-21
add-6 to both sides
3x+6-6=-21-6
3x=-27
divide by 3
3x/3 = -27/3
x=-9
the next odd integer will be x+2= -9+2=-7
The third will be x+4 = -9+4 = -5
..
The numbers are -5, -7, -9
And their sum = -21